# QADM-Net: Multi-Level Quality-Adaptive Dynamic Network for Reliable Multimodal Classification

### v2

@@ Abstract
Multimodal machine learning has achieved remarkable progress in many scenarios, but its reliability is undermined by varying sample quality. In this paper, we find that current multimodal classification methods lack dynamic networks for sample-specific depth and parameters to achieve reliable inference. To this end, a novel framework for multimodal reliable classification termed Multi-Level Quality-Adaptive Dynamic Multimodal Network (QADM-Net) is proposed. QADM-Net first adopts a novel approach based on noise-free prototypes and a classifier-free design to reliably estimate the quality of each sample at both modality and feature levels. It then achieves sample-specific network depth via the Global Confidence Normalized Depth (GCND) mechanism. By normalizing depth across modalities and samples, GCND effectively mitigates the impact of challenging modality inputs on dynamic depth reliability. Furthermore, QADM-Net provides sample-adaptive network parameters via the Layer-wise Greedy Parameter (LGP) mechanism driven by feature-level quality. The cross-modality layer-wise greedy strategy in LGP designs a reliable parameter prediction paradigm for multimodal networks with variable depths for the first time. Experiments conducted on four datasets demonstrate that QADM-Net significantly outperforms state-of-the-art methods in classification performance and reliability, exhibiting strong adaptability to data with diverse quality.
1 Introduction
@@
Figure 1: Empirical studies under varying data quality. We simulate data quality degradation by adding Gaussian noise, with $\sigma$ representing the noise intensity. (a) The classification accuracy corresponds to different network depths on the noisy modality with varying noise intensity. The results of three state-of-the-art reliable multimodal classification methods including CALM Zhou et al. (2023), GCFANet Zheng et al. (2024), and PDF Cao et al. (2024) are shown. (b) Visualization of the model parameters required to map samples with different features to their corresponding class centers. (c) The confidence estimation results on the same test samples provided by TCP Han et al. (2022) as a representative example after training on data with different noise intensities. More details are included in the supplementary material.
@@
In recent years, the accessibility of multimodal data has significantly enhanced prediction across various application scenarios Dou et al. (2022); Li et al. (2022); Sun et al. (2022); Elizalde et al. (2023). However, multimodal samples often display complex quality variations due to inherent differences in informativeness across modalities and the noise introduced by varying sensor conditions or environmental factors Zhang et al. (2024). This phenomenon poses substantial challenges for reliable prediction, especially in safety-critical domains such as computer-aided diagnosis and autonomous driving Feng et al. (2018); Nair et al. (2020); Han et al. (2022); Zhou et al. (2023). Consequently, extensive research Geng et al. (2021); Han et al. (2022); Zou et al. (2023); Zheng et al. (2023); Tellamekala et al. (2023); Zhou et al. (2023); Zhang et al. (2023); Zheng et al. (2024); Cao et al. (2024) has focused on developing reliable multimodal classification methods to handle data with varying quality. These methods evaluate the quality of samples and remove uninformative features in low-quality data, thus enhancing the reliability of learning results.
@@
However, empirical studies reveal several limitations of current reliable multimodal classification methods when dealing with data of varying quality. (1) Lack of network dynamics. We analyze the quality variation of multimodal samples at both the modality and feature levels. At the modality level, the optimal network depth grows as the modality quality declines (see the two curves for the same model in Figure 1 (a)). At the feature level, mapping feature vectors with different feature values to the most reliable representations (i.e., class centers) requires distinct network parameters, as shown in Figure 1 (b). Unfortunately, existing methods are static networks with fixed network depth and parameters. (2) Lack of reliable confidence estimation. These methods typically produce confidence estimation results that are sensitive to training noise, as shown in Figure 1 (c) and 5, leading to decreased reliability. Moreover, they cannot simultaneously perform multi-level (the modality and feature levels) quality estimation of a sample. Thus, we pose two key questions: (1) How can we design a robust multi-level quality estimation method to support reliable network adjustments? (2) How can we adjust the network’s depth and parameters to ensure reliability for a given multimodal sample?
@@
This paper addresses the above questions by proposing Multi-Level Quality-Adaptive Dynamic Multimodal Network (QADM-Net), a novel framework for reliable multimodal classification. QADM-Net first adopts noise-free prototype confidence estimation (NFCE) to reliably assess sample quality at both the modality and feature levels. Then, QADM-Net sequentially employs the Global Confidence Normalized Depth (GCND) and Layer-wise Greedy Parameter (LGP) mechanisms to adjust its depth and parameters based on modality- and feature-level qualities of the sample, respectively. Unlike existing confidence estimation methods, NFCE derives class probability from pre-optimized noise-free prototypes (i.e., noise-removed class centers) and a classifier-free design, instead of using network structures or classifiers. This design enables multi-level and reliable quality estimation by avoiding overfitting to training noise in the network structure and classifier. The GCND mechanism maps sample modality quality to network depth. Notably, it reduces the impact of extremely difficult modality inputs by normalizing depths across modalities of all samples, thereby enhancing the reliability of dynamic depth. The LGP mechanism further performs cross-modality parameters prediction aiming to maximize feature-level quality enhancement at each layer. This mechanism introduces a reliable parameter prediction framework across multimodal networks of varying depths for the first time.
@@
  • We delve into reliable multimodal learning from a novel perspective by portraying the dynamics of network depth and parameters to accommodate quality variations of different data at the modality and feature levels.
@@
  • The proposed confidence estimation method based on noise-free prototypes and classifier-free design surpasses existing methods in avoiding noise overfitting and achieving multi-level quality assessment.
@@
  • The two proposed dynamic mechanisms with cross-modality depth normalization and parameters prediction, design a new paradigm that endows dynamic networks with multimodal reliability for the first time.
  • The proposed QADM-Net outperforms state-of-the-art reliable multimodal learning methods in multiple benchmarks by a large margin and exhibits strong adaptability to data with various qualities.
Comparison with existing works.
We discuss the key differences between QADM-Net and other works. (1) Reliable multimodal classification methods: existing methods based on static networks, while QADM-Net possesses network dynamics and adaptability towards different input samples. (2) Dynamic neural networks (DNNs): Our method endows DNNs with multimodal reliability for the first time. Most DNNs, except for DynMM Xue and Marculescu (2023), are unimodal and cannot be extended to multimodal scenarios. All DNNs, including DynMM, commonly lack reliability in both confidence estimation and dynamic mechanism design. QADM-Net complements these methods by offering reliable multi-level confidence estimation and robust dynamic mechanisms based on sample’s quality and cross-modality relationships. (3) Confidence estimation methods: The proposed NFCE is orthogonal to existing confidence estimation methods, i.e., NFCE can be used to improve the reliability of these methods and achieve multi-level quality estimation.
2 Related Works
Due to space limitations, this section highlights only a selection of related works. For a more comprehensive overview, please refer to the supplementary material.
2.1 Reliable Multimodal Classification
Many studies have developed quality-aware methods to provide reliable multimodal classification results from low-quality multimodal data. Han et al. Han et al. (2022) modeled informativeness at the feature and modality levels, achieving trustworthy multi-modal feature fusion. Zhang et al. Zhang et al. (2023), Cao et al. Cao et al. (2024) achieved more robust and generalized multimodal fusion by dynamically assigning weights to each modality based on uncertainty estimation. Zheng et al. Zheng et al. (2023) achieved trustworthy multimodal classification via integrating feature and label-level confidence.
2.2 Confidence Estimation
In reliable multimodal classification methods, confidence estimation measures the quality of input features, serving as a necessary foundation for the model’s reliability. Some methods use Maximum-Class Probability (MCP) for confidence estimation Zou et al. (2023), while others adopt True Class Probability (TCP) Han et al. (2022); Cao et al. (2024), and some rely on energy scores Zhang et al. (2023). These methods commonly estimate confidence based on the class probability of input obtained by classifiers or network structures.
@@ 2.2 Confidence Estimation
Figure 2: The framework of the proposed QADM-Net. For an input sample $x_{i}$ from dataset $\mathcal{D}$: (a) the modality-level and feature-level quality of $x_{i}$ is estimated via Noise-free Prototype Confidence Estimation (NFCE). (b) The reliable depth of QADM-Net for $x_{i}$ is adjusted by Global Confidence Normalized Depth (GCND) based on modality-level quality. (c) The reliable parameters of QADM-Net for $x_{i}$ are adjusted by Layer-wise Greedy Parameter (LGP) based on feature-level quality.
2.3 Dynamic Neural Network
Designing dynamic mechanisms is a promising approach to promoting neural networks’ accuracy and computational efficiency Veit and Belongie (2018); Hua et al. (2019); Su et al. (2019); Han et al. (2021). Huang et al. (2017) utilized early exit to enable efficient inference for samples. Lin et al. Lin et al. (2023) achieved accuracy-speed trade-offs by applying different network routines to each sample. Xue et al. Xue and Marculescu (2023) introduced DynMM, a dynamic multimodal fusion approach that reduces computational costs while maintaining accuracy.
3 Proposed QADM-Net
@@
In this section, we provide a detailed illustration of the proposed QADM-Net. Given a multimodal dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$ containing $N$ samples with $M$ modalities belonging to $C$ classes. Multimodal classification methods aim to learn a neural network that maps every input sample $x_{i}=\{x_{i}^{m}\in\mathbb{R}^{d^{m}}\}_{i=m}^{M}$ to its corresponding class label $y_{i}\in\mathbb{R}^{C}$, where $d^{m}$ is the dimension of the $m$-th modality. In this work, we propose QADM-Net to provide reliable classification result $\hat{y}_{i}$ for input sample $x_{i}$ via sample-adaptive network depth and parameters.
3.1 Backbone of QADM-Net
We first introduce the hierarchical backbone structure of QADM-Net. QADM-Net consists of $M$ unimodal networks $N^{m}(m\in[1,M])$. Each $N^{m}$ comprises $\textbf{D}^{m}$ network blocks $B^{m,t}(t\in[1,\textbf{D}^{m}])$. Each network block $B^{m,t}$ contains a fully connected layer with parameters $\theta^{m,t}=[W^{m,t},b^{m,t}]$, where $W^{m,t},b^{m,t}$ refers to weight and bias.
@@ 3.1 Backbone of QADM-Net
Given an input sample $x_{i}=\{x_{i}^{m}\}_{m=1}^{M}$, the depth $\textbf{D}^{m}$ of each unimodal network is adjusted to $\textbf{D}^{m}_{i}$, and all network block parameters $\theta^{m,t}$ are predicted and assigned as $\theta^{m,t}_{i}=[W^{m,t}_{i},b^{m,t}_{i}]$, as elaborated in Section 3.3 and 3.4. Input feature vectors of each modality are gradually enhanced through network blocks in their respective unimodal networks to obtain high-quality output features $f_{i}=\{f_{i}^{m}\}_{m=1}^{M}$:
$f_{i}^{m}$ $=N^{m}(x_{i}^{m})=B^{m,\textbf{D}_{i}^{m}}\circ...\circ B^{m,1}(x_{i}^{m}),$ (1)
where $\circ$ refers to the composition of two operations. Let $I^{m,t}_{i}$ and $I^{m,t+1}_{i}$ denote the input and output of the network block $B^{m,t}$. In particular, $I^{m,1}_{i}=x_{i}^{m},I^{m,\textbf{D}_{i}^{m}+1}_{i}=f_{i}^{m}$. The quality enhancement by each network block $B^{m,t}:I^{m,t+1}_{i}=B^{m,t}(I^{m,t}_{i})$ can be formulated as:
@@
$\sigma$ is the sigmoid function. $q_{i}^{m,t}\in\mathbb{R}^{d^{m}}$ is the feature-level informativeness vector of $I_{i}^{m,t}$ learned by $B^{m,t}$, with each element reflects the informativeness of each feature value in $I^{m,t}_{i}$. The quality enhancement is achieved by retaining informative features while removing redundant ones through element-wise multiplication $\odot$ between $q^{m,t}_{i}$ and $I^{m,t}_{i}$.
@@
3.2 Noise-free Prototype Confidence Estimation

### v3

# Multi-QuAD: Multi-Level Quality-Adaptive Dynamic Network for Reliable Multimodal Classification

@@ Abstract
Multimodal machine learning has achieved remarkable progress in many scenarios, but its reliability is undermined by varying sample quality. This paper finds that existing reliable multimodal classification methods not only fail to provide robust estimation of data quality, but also lack dynamic networks for sample-specific depth and parameters to achieve reliable inference. To this end, a novel framework for multimodal reliable classification termed Multi-level Quality-Adaptive Dynamic multimodal network (Multi-QuAD) is proposed. Multi-QuAD first adopts a novel approach based on noise-free prototypes and a classifier-free design to reliably estimate the quality of each sample at both modality and feature levels. It then achieves sample-specific network depth via the Global Confidence Normalized Depth (GCND) mechanism. By normalizing depth across modalities and samples, GCND effectively mitigates the impact of challenging modality inputs on dynamic depth reliability. Furthermore, Multi-QuAD provides sample-adaptive network parameters via the Layer-wise Greedy Parameter (LGP) mechanism driven by feature-level quality. The cross-modality layer-wise greedy strategy in LGP designs a reliable parameter prediction paradigm for multimodal networks with variable architecture for the first time. Experiments conducted on four datasets demonstrate that Multi-QuAD significantly outperforms state-of-the-art methods in classification performance and reliability, exhibiting strong adaptability to data with diverse quality.
Index Terms:
Multimodal learning, classification, reliability, trustworthy, multimodal noise, dynamic network.
I Introduction
@@
Figure 1: Empirical studies under varying data quality. We simulate data quality degradation by adding Gaussian noise to one of the modalities on BRCA dataset, with $\sigma$ representing the noise intensity. (a) Classification accuracy (ACC) of three state-of-the-art methods (GCFANet [1], PDF [2], CALM [3]) under different depths of their unimodal network depths corresponding to the noisy modality. (b) Visualization of the model parameters required to map samples with different features to their corresponding class centers. (c) The confidence estimation results on the same test samples provided by TCP [4] as a representative example after training on data with different noise intensities. More details of the observation experiments are included in the Supplemental Materials.
@@
The accessibility of multimodal data has significantly enhanced the performance of machine learning in various scenarios [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16], especially in multimedia applications [17, 18, 19]. However, multimodal samples often exhibit complex quality variations, which are primarily attributed to the inherent differences in the informativeness across modalities [20, 21, 22] and the noise introduced by varying sensor conditions or environmental factors [23, 24, 25]. This phenomenon poses substantial challenges for reliable model prediction, especially in safety-critical domains such as computer-aided diagnosis and autonomous driving [26, 27, 4, 3]. To tackle this challenge, extensive research has explored the variability in multimodal sample quality at both modality and feature levels [24, 4, 28, 29, 30, 3, 31, 1, 2], significantly enhancing classification reliability by removing low-quality and uninformative features. The modality level refers to the quality variations of modalities among samples, while the feature level pertains to the quality differences of individual feature values within each modality’s feature vector across samples.
@@
However, empirical studies reveal several limitations of current reliable multimodal classification methods when dealing with data of varying quality. (1) Lack of network dynamics. We introduced Gaussian noise to one modality of the data to simulate modality-level quality degradation and implemented versions of the state-of-the-art model with different combinations of unimodal network depths. Fig. 1a shows that as the quality degrades (increasing noise intensity $\sigma$ from 1 to 5), the unimodal networks corresponding to the affected modality in all SOTA models consistently require greater depth to achieve the optimal accuracy. Additionally, the experimental results shown in Fig. 1b indicate that different network parameters are required to map feature vectors of samples with inherent feature-level quality differences to the most reliable representation (e.g., their corresponding class centers). Unfortunately, existing methods are static networks with fixed network depth and parameters. (2) Lack of reliable confidence estimation. These methods typically provide confidence estimation results that overfit the training noise, leading to decreased reliability. As shown in Fig. 1c and 6, existing confidence estimation methods yield significantly different results on the same test samples after being trained on data with different noise settings. Moreover, they cannot simultaneously perform quality estimation of a sample at both modality and feature levels. Thus, we pose two key questions: (1) How can we design a robust multi-level quality estimation method to support reliable network adjustments? (2) How can we adjust the network’s depth and parameters to ensure reliability for a given multimodal sample?
@@
This paper addresses the above questions by proposing Multi-level Quality-Adaptive Dynamic multimodal network (Multi-QuAD), a novel framework for reliable multimodal classification. Multi-QuAD first adopts noise-free prototype confidence estimation (NFCE) to reliably estimate sample quality at both the modality and feature levels. Then, Multi-QuAD sequentially employs the Global Confidence Normalized Depth (GCND) and Layer-wise Greedy Parameter (LGP) mechanisms to adjust its depth and parameters based on modality- and feature-level qualities of the sample, respectively. Unlike existing confidence estimation methods, NFCE derives class probability via a classifier-free design using noise-free prototypes (noise-removed class centers) pre-optimized based on inter-modality and inter-class relationships, instead of using network structures or classifiers. This design not only enables multi-level quality estimation but also facilitates reliable quality estimation by avoiding overfitting to training noise in the network structure or classifier. The GCND mechanism maps sample modality quality to network depth. Notably, it reduces the impact of extremely difficult modality inputs by normalizing depths across modalities of all samples, thereby enhancing the reliability of dynamic depth. The LGP mechanism further performs cross-modality parameters prediction driven by a novel shallow-to-depth layer-wise greedy strategy aiming to maximize feature-level quality enhancement at each network layer. This mechanism introduces an efficient and reliable parameter prediction framework on a flexible multimodal framework with dynamic depths for the first time.
@@
  • We propose Multi-QuAD, which introduces a novel perspective on reliable multimodal classification by dynamically adapting network depth and parameters to varying data quality at both modality and feature levels. Multi-QuAD significantly outperforms state-of-the-art methods in accuracy and reliability across multiple benchmarks.
@@
  • The proposed classifier-free method NFCE leverages noise-free prototypes pre-optimized by inter-modality and inter-class relations, outperforming existing confidence estimation methods in avoiding noise overfitting and enabling multi-level quality assessment.
@@
  • The two proposed dynamic mechanisms, GCND and LGP, with cross-modality depth normalization and layer-wise greedy parameter prediction, design a new paradigm that endows dynamic networks with multimodal reliability for the first time.
II Related Works
This section reviews some related works on reliable multimodal classification, classification confidence estimation, and dynamic neural networks.
II-A Reliable Multimodal Classification
Multimodal learning has achieved significant success in enhancing model performance across numerous application fields [32, 33, 34, 35, 36, 37, 38, 39]. However, multimodal data often contains noise, leading to the emergence of low-quality data, which in turn affects the reliability of multimodal learning. To this end, many studies have developed quality-aware methods to provide reliable multimodal classification results from low-quality multimodal data. Some methods have focused on addressing modality-level quality issues. For example, Han et al. [40] have parameterized the evidence of different modality features using Dirichlet distribution and fused each modality based on their evidence using Dempster-Shafer theory. Geng et al. [24] have proposed the DUA-Nets, which achieved multimodal representation learning based on modality uncertainty. Zhang et al. [31], Cao et al. [2] have achieved more robust and generalized multimodal fusion by dynamically assigning weights to each modality based on their uncertainty. In contrast, some studies address data quality issues across multiple levels. For instance, Han et al. [4] have modeled informativeness at the feature and modality levels, achieving trustworthy multimodal feature fusion. Zhou et al. [3] have learned trustworthy features in each modality and achieved confidence-aware fusion among modalities. Despite their effectiveness, these methods are all based on static networks, whose limitations have been analyzed in Sec. I. To address this, we propose Multi-QuAD, which possesses network dynamics and adaptability towards different input samples.
II-B Confidence Estimation
In reliable multimodal classification methods, confidence estimation can reflect the quality of the model input, serving as a necessary foundation for the model’s reliability. In the early years, Bayesian neural networks [41, 42, 43], predictive confidence [44], and Dempster-Shafer theory [40] were proposed for confidence estimation. Predictive confidence measures the consistency between the model’s predicted class probabilities and its empirical accuracy, which typically refers to the accuracy in classification tasks. Dempster-Shafer theory (DST) measures uncertainty by assigning probabilities to sets rather than individual events, allowing for handling information at different precision levels without introducing additional assumptions. In recent years, several easily implementable, lightweight confidence estimation methods have been proposed. For example, Maximum-Class Probability (MCP) [28] uses the maximum value of the classification probability predicted by the model as the confidence assessment. However, MCP suffers from significant overconfidence issues [45, 4]. To address this, methods such as energy score [31] and True-Class-Probability (TCP) [4, 2] have been proposed, which leverage the model’s predicted probability to obtain more reliable confidence assessments. However, these methods, which utilize classification probabilities predicted by network architectures or classifiers, suffer from overfitting to training noise and are incapable of multi-level quality assessment. To address these common issues, we propose noise-free prototype confidence estimation (NFCE) based on noise-free prototypes and a classifier-free design.
@@ II-B Confidence Estimation
Figure 2: The framework of the proposed Multi-QuAD (better viewed in colour). Without loss of generality, this figure illustrates the case of two modalities, with blue and yellow representing different modalities. For an input sample $x_{i}$ from dataset $\mathcal{D}$: (a) the modality-level and feature-level quality of $x_{i}$ is estimated via Noise-free Prototype Confidence Estimation (NFCE). (b) The reliable depth of Multi-QuAD for $x_{i}$ is adjusted by Global Confidence Normalized Depth (GCND) based on modality-level quality. (c) The reliable parameters of Multi-QuAD for $x_{i}$ are adjusted by Layer-wise Greedy Parameter (LGP) driven by feature-level quality. The detailed implementation of $\textbf{LGP}^{t}$ at each layer will be demonstrated in Fig. 3.
II-C Dynamic Neural Network
Designing dynamic architectures or parameters is a promising approach to promoting neural networks’ accuracy, computational efficiency, and adaptability in multiple tasks [46, 47, 48, 49]. In image classification tasks, several works [50, 51, 52] have utilized the concept of early exit to enable efficient inference for samples. In natural language processing tasks, many studies [53, 54] have effectively improved inference speed and addressed the overthinking issue by allowing samples to be output at different network layers. In object detection tasks, Lin et al. [55] have achieved accuracy-speed trade-offs by applying different network routines to different samples. In the multimodal task, DynMM [56] has employed a dynamic fusion strategy to achieve high computational efficiency in emotion recognition. However, dynamic neural networks (DNNs) generally lack multimodal reliability. Most DNNs, except for DynMM [56], are unimodal and cannot be extended to multimodal scenarios. All DNNs, including DynMM, commonly lack reliability in both confidence estimation and dynamic mechanism design. The proposed Multi-QuAD complements these methods by offering reliable multi-level confidence estimation and robust dynamic mechanisms based on the sample’s quality and cross-modality relationships, which endows DNNs with multimodal reliability for the first time.
III Proposed Approach
@@
In this section, we provide a detailed illustration of the proposed Multi-QuAD. Given a multimodal dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$ containing $N$ samples with $M$ modalities belonging to $C$ classes. Multimodal classification methods aim to learn a neural network that maps every input sample $x_{i}=\{x_{i}^{m}\in\mathbb{R}^{d^{m}}\}_{i=m}^{M}$ to its corresponding class label $y_{i}\in\mathbb{R}^{C}$, where $d^{m}$ is the dimension of the $m$-th modality. In this work, we propose Multi-QuAD to provide reliable classification result $\hat{y}_{i}$ for input sample $x_{i}$ via sample-adaptive network depth and parameters.
III-A Backbone of Multi-QuAD
@@
We first introduce the hierarchical backbone structure of Multi-QuAD. Multi-QuAD is an adaptive network that dynamically adjusts according to different input samples. It consists of $M$ unimodal networks, each formed by stacking an adjustable number of network blocks. Each block is composed of a single neural network layer, with the parameters dynamically predicted based on the input sample. Given an input sample $x_{i}=\{x_{i}^{m}\}_{m=1}^{M}$, the adjusted Multi-QuAD is denoted as $QUAD_{i}$, where each unimodal network is denoted as $N_{i}^{m}(m\in[1,M])$. The depth of each unimodal network $N_{i}^{m}$ is adjusted to $\textbf{D}_{i}^{m}$ based on the input $x_{i}$, i.e., stacking $\textbf{D}_{i}^{m}$ network blocks in $N_{i}^{m}$, as elaborated in Sec. III-C. The $t$-th $(t\in[1,\textbf{D}_{i}^{m}])$ network block in $N_{i}^{m}$ is denoted as $B_{i}^{m,t}$, and its network layer parameters are denoted as $\theta_{i}^{m,t}$. $\theta_{i}^{m,t}=[W_{i}^{m,t},b_{i}^{m,t}]$ where $W_{i}^{m,t}$ and $b_{i}^{m,t}$ represent the weight and bias, respectively. The parameters $\theta_{i}^{m,t}$ of each block are dynamically predicted during the inference process for the sample $x_{i}$, as elaborated in Sec. III-D. Input feature vectors of each modality in $x_{i}$ are gradually enhanced through network blocks in their respective unimodal networks, and finally obtain high-quality output features $f_{i}=\{f_{i}^{m}\}_{m=1}^{M}$:
$f_{i}^{m}$ $=N^{m}_{i}(x_{i}^{m})=B^{m,\textbf{D}_{i}^{m}}_{i}\circ...\circ B^{m,1}_{i}(x_{i}^{m}),$ (1)
where $\circ$ refers to the composition of two operations. The calculation in the network block $B^{m,t}_{i}$ can be formulate as:
@@
$I^{m,t}_{i}$ and $I^{m,t+1}_{i}$ denote the input and output of the network block $B^{m,t}_{i}$, respectively, and $I^{m,1}_{i}=x_{i}^{m},I^{m,\textbf{D}_{i}^{m}+1}_{i}=f_{i}^{m}$. $\sigma$ is the sigmoid function. $q_{i}^{m,t}\in\mathbb{R}^{d^{m}}$ is the feature-level informativeness vector of $I_{i}^{m,t}$ learned by $B^{m,t}_{i}$, with each element reflects the informativeness of each feature value in $I^{m,t}_{i}$. The quality enhancement is achieved by retaining informative features while removing redundant ones through element-wise multiplication $\odot$ between $q^{m,t}_{i}$ and $I^{m,t}_{i}$.
@@
Figure 3: The detailed implementation of cross-modality greedy parameter prediction at the $t$-th layer ($\textbf{LGP}^{t}$).
III-B Noise-free Prototype Confidence Estimation